HETAV PATEL

Grade 11,

in an intrinsic semiconductor the band gap is 1.2 then the ratio of number of charge carriers at 600k and 300k

in an intrinsic semiconductor the band gap is 1.2 then the ratio of number of charge carriers at 600k and 300k

Grade:8

1 Answers

Vikas TU
6 years ago
 Here the band hole = 1.2 
Temperature (T1) = 600 K 
Temperature (T2) = 300k 
As we ol realize that 
N12/Ni1= (n0[ exp(- Eg/2KB600)])/( n0[ exp(- Eg/2KB300)]) 
= exp Eg/2KB[1/300-1/600] 
= exp (1.2/2*8.62* 10^-5 * 1/600) 
= exp[11.6] 
= 1.09 * 10^5